OpenStudy (anonymous):

One number exceeds another number by 12. The sums of the numbers is 80. What are the numbers?

OpenStudy (anonymous):

Take the two numbers to be x and x+12, then add them

hero (hero):

x + y = 80 x = y + 12 y = 80 - x y = x - 12 y = y 80 - x = x - 12 x + x = 80 + 12 2x = 92 x = 46 y = 34